OpenStudy (anonymous):

if k= 2pi/45 what is the period of y=coskx a)45/2 b)45 c)45/2pi d)2pi/45

OpenStudy (anonymous):

period of \(\cos(kx)\) is \(\frac{2\pi}{k}\)

OpenStudy (anonymous):

\[\frac{2\pi}{\frac{2\pi}{45}}=\frac{2\pi\times 45}{2\pi}=45\]